OpenStudy (lgbasallote):

\[\begin{array}l\color{red}{\text{L}}\color{orange}{\text{G}}\color{#9c9a2e}{\text{B}}\color{green}{\text{A}}\color{blue}{\text{R}}\color{purple}{\text{I}}\color{purple}{\text{D}}\color{red}{\text{D}}\color{orange}{\text{L}}\color{#9c9a2e}{\text{E}}\color{green}{\text{!}}\color{blue}{\text{!}}\color{purple}{\text{!}}\color{purple}{\text{}}\end{array}\] 965 = 21 578 = 51 794 = ???

OpenStudy (anonymous):

31..

OpenStudy (anonymous):

answer is 41

OpenStudy (anonymous):

+ 5 = 36..

OpenStudy (lgbasallote):

no and no

OpenStudy (lgbasallote):

and no

mathslover (mathslover):

44 ?

OpenStudy (anonymous):

Is there any logic in this ??

OpenStudy (lgbasallote):

it sure isn't random

mathslover (mathslover):

yes it is of reasoning or mental ability..

OpenStudy (anonymous):

what is this

OpenStudy (lgbasallote):

this is one of the notoriously absurd lgbariddles

mathslover (mathslover):

:P wait

OpenStudy (lgbasallote):

there's actually a pattern

mathslover (mathslover):

yes..

OpenStudy (anonymous):

but i havent see any pattern

mathslover (mathslover):

is it 81 lgba?

OpenStudy (lgbasallote):

that's why it is an lgbariddle

OpenStudy (lgbasallote):

no @mathslover

mathslover (mathslover):

thanks

OpenStudy (anonymous):

all i see is that the patterns are paired from each side-same colour.but how do introduce numbers

OpenStudy (lgbasallote):

i just noticed mathslover is almost green

OpenStudy (anonymous):

Almost..

mathslover (mathslover):

"almost" I am not going to be soon since I have half yearly exams from ten

OpenStudy (lgbasallote):

mathslover will turn green when i become 100

mathslover (mathslover):

:P it smells true to me

OpenStudy (anonymous):

Is the answer 75 ??

OpenStudy (lgbasallote):

still no

OpenStudy (lgbasallote):

do you even have bases for those numbers?

OpenStudy (anonymous):

I will guess it again..

OpenStudy (anonymous):

another guess by waterineyes

OpenStudy (lgbasallote):

sure. you just have 999,999,999, 999,999,999, 999,999,999, 999,999,999, 999,999,999, 999,999,999, 999,999,999, 999,999,999, 999,999,999, 999,999,999, numbers to choose from

mathslover (mathslover):

NFE = BA EGH = EA GID = __ "Hint" tips hat :P

OpenStudy (lgbasallote):

interesting

mathslover (mathslover):

I know the answer but will not declare here... let me see who is the other one who knows

OpenStudy (lgbasallote):

as you please

OpenStudy (anonymous):

Don't worry lgba number is less than 100 so I have 99 numbers to choose from..

OpenStudy (anonymous):

So 96 guesses are still left.. ha ha ha..

OpenStudy (lgbasallote):

94 actually

OpenStudy (anonymous):

I am counting my only..

OpenStudy (lgbasallote):

as you please

OpenStudy (lgbasallote):

seems a lot of people are quitting....

OpenStudy (anonymous):

is there a pattern behind it?

OpenStudy (lgbasallote):

yes. there is.

OpenStudy (lgbasallote):

just like all my other lgbariddles

OpenStudy (anonymous):

Is it 73 ??

OpenStudy (lgbasallote):

still no. guess again

OpenStudy (anonymous):

13, 23, 33, 43, 53, 63, 83, 93 any of these ??

OpenStudy (lgbasallote):

lol

OpenStudy (lgbasallote):

just look for the pattern

OpenStudy (anonymous):

Oh then we got it guys answer is within those numbers..

OpenStudy (anonymous):

hmm..im seeing.. 9+6+5 = 20 5+7+8 = 20 7+9+4 = 20 What may the pattern be

OpenStudy (lgbasallote):

hmm nice....

mathslover (mathslover):

that is what I said but that is not any good pattern being shown for our answer

mathslover (mathslover):

+1 +31 +61 is the answer 81 ? I will ask again

OpenStudy (lgbasallote):

it's not more right than when you first asked

hartnn (hartnn):

my first guess 12 it that right ?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

Ha ha ha ha

OpenStudy (lgbasallote):

why would you say 12?

OpenStudy (anonymous):

Oh so you may be right hartnn

OpenStudy (lgbasallote):

i think he's on to a pattern

hartnn (hartnn):

since u have given very less pairs to deal with,m ignoring first 9,5(MSD) 5+3=8 so 2+3=5 thats y 1 (5-4=1) and 7-6=1 so 9-7=2 ......this logic may be wrong,but is 12 correct?

OpenStudy (lgbasallote):

nope

OpenStudy (anonymous):

hint please

OpenStudy (lgbasallote):

PEMDAS

OpenStudy (anonymous):

whats that

OpenStudy (anonymous):

Yeah I am still doing that only @lgbasallote

OpenStudy (lgbasallote):

Parenthesis Exponent Multiplication Division Addition Subtraction

OpenStudy (anonymous):

Tell me if we use division here or not ??

OpenStudy (lgbasallote):

okay...a more subtle hint the numbers on the LHS dont make a pattern the numbers on the RHS dont make a pattern

OpenStudy (anonymous):

Oh wow I got your second hint.. Can we move to third hint now @lgbasallote

OpenStudy (anonymous):

does not make sense; where is the pattern then

OpenStudy (lgbasallote):

the equation is the pattern

OpenStudy (anonymous):

In this sign " = " because this is what left then @Omniscience

OpenStudy (anonymous):

so there is a pattern in this \(965 = 21\) and the rest does not not matter?

OpenStudy (lgbasallote):

if you can find out why 965 = 21 then you can find out how to get 794 the second equation is merely something you can test your theories on

OpenStudy (anonymous):

ok; whats parenthesis?

OpenStudy (anonymous):

brackets?

OpenStudy (anonymous):

and whats exponent lol

OpenStudy (lgbasallote):

i didn't say you'll use all of PEMDAS

OpenStudy (lgbasallote):

but it does follow PEMDAS

OpenStudy (anonymous):

division is being used here ??

OpenStudy (anonymous):

utterly ambiguous. 9 6 5 => 9*6*5 = > 270 => 270/ (3) => 90 + (9+6+5) => 110 what is this i dont get it; can you reuse the numbers?

OpenStudy (lgbasallote):

you can...

OpenStudy (anonymous):

am i on the right track?

ganeshie8 (ganeshie8):

29

OpenStudy (anonymous):

That is what I was just writing @ganeshie8 ha ha ha..

ganeshie8 (ganeshie8):

lol the other riddle lgba posted in physics group is interesting

OpenStudy (anonymous):

link

mathslover (mathslover):

I did search on net but no link :P

mathslover (mathslover):

@lgbasallote hint pls

OpenStudy (unklerhaukus):

??? = 29

mathslover (mathslover):

how ??????

OpenStudy (unklerhaukus):

-,+,*

mathslover (mathslover):

-,+, * how? I didn't get you

OpenStudy (unklerhaukus):

\[-9+6\times5 = 21\] \[-5+7\times8 = 51\] \[-7+9\times4 = 29\]

mathslover (mathslover):

wowowowowwowowow ... you are right...and this can not be neglected...